Jeric is standing outside the Physics laboratory 7m from the wall. Michael
John, by a window 5m above the ground, tosses a ball horizontally. Find the
speed that Michael John should give the ball for it to reach jeric
Can you show me the conplete solution than you!
1
Expert's answer
2020-12-03T06:59:57-0500

The key thing here is that the horizontal distance passed by the ball (7 m) is defined by the horizontal velocity and time of flight, while the time of flight is defined by height. Begin with the last:


"t=\sqrt{2h/g}=\sqrt{\frac{2\cdot5}{9.8}}=1.01\text{ s}."


The range of 7 m is


"R=vt, \\\space\\
v=\frac Rt=\frac{7}{1.01}=6.93\text{ m/s}."